# Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness

### v1

@@ Abstract
Deep learning models achieve strong performance across various domains but often rely on spurious correlations, making them vulnerable to distribution shifts. This issue is particularly severe in subpopulation shift scenarios, where models struggle in underrepresented groups. While existing methods have made progress in mitigating this issue, their performance gains are still constrained. They lack a rigorous theoretical framework connecting the embedding space representations with worst-group error. To address this limitation, we propose Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness (SCER), a novel approach that directly regularizes feature representations to suppress spurious cues. We show theoretically that worst-group error is influenced by how strongly the classifier relies on spurious versus core directions, identified from differences in group-wise mean embeddings across domains and classes. By imposing theoretical constraints at the embedding level, SCER encourages models to focus on core features while reducing sensitivity to spurious patterns. Through systematic evaluation on multiple vision and language, we show that SCER outperforms prior state-of-the-art studies in worst-group accuracy. Our code is available at https://github.com/MLAI-Yonsei/SCER.
@@ 1 Introduction
Deep neural architectures have transformed performance standards in many domains. However, they are still highly vulnerable to spurious correlations in real-world datasets (yang2023mitigating; izmailov2022feature; sagawa2020investigation). Spurious correlations refer to patterns in the training data that are statistically associated with target labels without a true causal relationship. Such correlations are often the result of biases in the dataset, causing models to fail when faced with distribution shifts (ye2024spurious). Recent studies have shown that this issue is particularly severe in subpopulation shift scenarios, where the distribution of certain subpopulations in the test data differs from the training distribution (sagawadistributionally; yang2023change). Models which are trained with Empirical Risk Minimization (ERM) tend to over-rely on these spurious correlations, as ERM minimizes the average loss without considering subpopulation imbalances (vapnik2013nature; izmailov2022feature). As a result, models rely on easily learned but non-robust features, leading to poor worst-group accuracy and unreliable performance on underrepresented subpopulations.
@@
Numerous approaches have been developed to handle this challenge, which can be broadly grouped into several categories (yang2023change). Subpopulation robustness methods  (sagawadistributionally; duchi2021learning; pmlr-v139-liu21f; yao2022improving; deng2023robust) focus on improving worst-group performance through various techniques such as reweighting, multi-stage training, or specialized loss functions. Domain invariant methods (arjovsky2019invariant; sun2016deep) attempt to learn features that generalize across different domains by aligning feature distributions. Data augmentation approaches (zhang2018mixup) create synthetic training examples to improve generalization. Class imbalance methods (japkowicz2000class; cui2019class; ren2020balanced) address the uneven distribution of samples across classes. However, these approaches primarily influence model learning indirectly through sample reweighting or feature alignment, without explicitly constraining how spurious features are encoded within the embedding space. Consequently, spurious correlations may persist, limiting robustness under distribution shifts.
@@
To address this limitation, we propose Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness (SCER), a novel method that directly regularizes feature representations to prevent reliance on spurious features. As illustrated in Figure 1, SCER operates at the embedding level, imposing explicit constraints that guide the model toward learning more robust features. In this study, we provide a new theoretical relationship between worst-group error and the structural properties of representation space under subpopulation shift scenarios, where spurious correlations exist between domain and label. We newly observe two key factors that guarantee the worst-group error from our theoretical analysis. First, for instance, sharing the same label but differing in domain, we regularize their embeddings to capture essential features for classification. Second, for instances sharing the same domain but with different labels, we introduce additional constraints to reduce the impact of domain-specific spurious cues. Through this approach, SCER prevents the model from overfitting to domain-level artifacts, enabling it to learn robust core features more effectively.
@@
By structuring the model’s representation space, SCER improves worst-group accuracy while maintaining strong overall performance. Through extensive experiments on both image and text datasets, we demonstrate that SCER outperforms prior methods in worst-group accuracy. We also provide quantitative and qualitative evidence that embedding-level constraints yield robust feature representations even under spurious domain shifts, confirming the effectiveness of the SCER.
Figure 1: Overview of the Spurious Correlation-Aware Embedding Regularization (SCER) framework. SCER distinguishes spurious and core features via group-wise embedding differences, then regularizes to minimize spurious loss and maximize core loss for domain bias robustness.
@@ 2.1 Subpopulation Shift
Machine learning models degrade under subpopulation shift, where test subpopulation distributions differ from training data (yang2023change). ERM’s focus on average training loss makes models susceptible to spurious correlations (vapnik2013nature; izmailov2022feature). To address this issue, several approaches have been proposed which can be broadly categorized as sample reweighting and data augmentation strategies. GroupDRO (sagawadistributionally) reduces worst-group error by dynamically upweighting samples from groups with higher loss. LfF (bahng2020learning) employs a two-stage approach that first detects spurious correlations by training a biased model and then debiases a second model by reweighting the loss gradient. GIC (han2024improving) improves the accuracy of group inference by leveraging spurious attribute associations. For data augmentation, LISA (yao2022improving) enhances robustness by interpolating between subpopulation samples to encourage invariant representations, while PDE (deng2023robust) gradually expands the training set, allowing models to first learn core features before incorporating more diverse samples. While prior research focuses primarily on sample reweighting or data augmentation, we propose SCER to directly regularize the embedding space and reduce reliance on spurious features. By explicitly penalizing dependence on spurious cues at the representation level, SCER significantly improves robustness across various distribution shift benchmarks.
@@ 2.2 Embedding Space
Recent studies in the text and image domains have actively explored methods to refine latent spaces, aiming to improve representation quality and better capture the underlying structure among samples (patil2023survey; mingexploit). However, within the domain generalization setting, research has primarily focused on learning invariant features rather than explicitly modifying the embedding space itself. Previous work has attempted to mitigate unintended correlations arising from background, lighting, or dataset-specific biases (yang2023change). hermann2020shapes analyzed how networks differentiate robust and spurious features, while LfF  (bahng2020learning) introduced a two-stage training approach to guide model learning. More recently, ElRep (wen2025elastic) applies norm penalties to final-layer representations with demonstrated practical effectiveness. However, these methods typically guide model learning indirectly rather than explicitly constraining spurious feature encoding in the embedding space (patil2023survey; mingexploit). In contrast to previous approaches, we propose to directly impose an explicit loss function on the embedding space, ensuring that learned representations minimize reliance on spurious correlations and improve generalization under distribution shifts. Our work addresses these theoretical gaps by providing rigorous justification for embedding regularization and establishing clear connections to robust generalization.
@@ 3.1 Problem Setting
To further enhance robustness, we introduce a regularization approach motivated by the geometric decomposition of worst-group error into spurious and core components within the embedding space. Rather than addressing group errors solely through loss optimization, we incorporate structural regularization that encourages the classifier to align with label-consistent Core Feature while penalizing dependence on domain-specific Spurious Feature. These regularization terms, formulated through directional correlations in the embedding space, are detailed in the following section.
@@ 3.2 SCER
We propose SCER, which enhances the generalization performance of classifiers by decomposing the worst-group error and applying embedding-level regularization to each component as shown in Figure 2. Let $x_{\text{emb}}=f_{w}(x)\in\mathbb{R}^{p}$ denote the embedding vectors of input $x$, obtained by a feature extractor $f_{w}$ with parameters $w$, where $p$ is the embedding dimension. For each subpopulation defined by a label–domain pair $(y,d)$, we define its mean embedding as
@@
We compute mean embeddings as representative vectors for each subpopulation, enabling decomposition of worst-group error into two primary components
Spurious Feature The embedding mean difference in feature representations across different domains within the same class. Since we assume that label–domain pairs $(y,d)$ may exhibit spurious correlations, such differences should be minimized to prevent the model from overfitting to domain-specific artifacts. Formally, the spurious difference for a given class $y$ is defined as
@@
Core Feature The embedding mean difference in feature representations between different classes within the same domain. This quantity captures the true signal that the classifier should leverage for robust prediction. Formally, the core difference for a given domain $d$ is defined as
@@
To ensure proper interpretation based on theoretical insights, we normalize these direction vectors using the $\Sigma$-norm, defined as $\|v\|_{\Sigma}=\sqrt{v^{\top}\Sigma v}$ for $v\in\mathbb{R}^{p}$, where $\Sigma\in\mathbb{R}^{p\times p}$ denotes the empirical covariance matrix of the embedding vectors. This normalization accounts for the geometric structure of the embedding space, producing scalar quantities Spurious Magnitude $\|\Delta_{\text{spur}}\|_{\Sigma}$ and Core Magnitude $\|\Delta_{\text{core}}\|_{\Sigma}$ for use in regularization.
We also derive regularization terms by analyzing how classifier weights align with the Spurious and Core Directions. The key idea is that overreliance on Spurious Directions leads to poor generalization on minority groups, while alignment with Core Directions supports robust performance. Let $\beta^{*}=[\beta^{*}_{1},\dots,\beta^{*}_{j},\dots,\beta^{*}_{m}]\in\mathbb{R}^{p\times m}$ denote the current weight matrix of the classifier $f_{\beta}$, where each column $\beta^{*}_{j}\in\mathbb{R}^{p}$ corresponds to the weight vector for class $j$. During training, we use the current classifier weights to compute the alignment measures. To quantify the alignment between classifier weights and directional signals, we compute the average class-wise correlation with each direction, using the $\Sigma$-norm . We refer to these measures as Weight-Spurious Alignment and Weight-Core Alignment
@@
These correlation terms measure how strongly the classifier’s decision boundaries align with spurious or core feature directions, forming the basis of our embedding-level regularization.
@@
The Spurious Loss $\mathcal{L}_{\text{spur}}$ penalizes alignment with spurious directions, scaled by the magnitude of spurious variation in the data. The Core Loss $\mathcal{L}_{\text{core}}$ encourages alignment with core directions, with the negative sign ensuring that stronger core alignment reduces the overall loss. Our loss formulation penalizes alignment with spurious directions and encourages alignment with core directions. By introducing control parameters $\lambda_{\text{spur}}$ and $\lambda_{\text{core}}$, we derive our embedding loss function:
@@ 3.3 Theoretical Analysis of SCER
We now present a theoretical analysis that formally decomposes the worst-group error. Detailed proofs and generalizations are provided in Appendix A.2.
@@
Consider a classification problem where each data point $x\in\mathcal{X}$ is associated with a label $\mathcal{Y}=\{y_{-1},y_{+1}\}$ and a domain $\mathcal{D}=\{d_{R},d_{G}\}$. Each pair $(y,d)\in\mathcal{Y}\times\mathcal{D}$ defines a subpopulation. We assume that the data follows a group-conditional Gaussian distribution $x\mid(y,d)\sim\mathcal{N}(\mu^{(y,d)},\Sigma)$ where $\Sigma$ is positive definite. We further assume that $\Delta_{\text{core}}=\mu^{(y_{+1},d)}-\mu^{(y_{-1},d)}$ is constant across $d$, $\Delta_{\text{spur}}=\mu^{(y,d_{R})}-\mu^{(y,d_{G})}$ is constant across $y$, $\Sigma^{(d_{R})}=\Sigma^{(d_{G})}=\Sigma$, $\mathbb{P}(y=\pm 1)=\mathbb{P}(d\in\{d_{R},d_{G}\})=\frac{1}{2}$, and $\mathbb{P}(y,d)\approx 0$ for some pairs $(y,d)$ under extreme spurious correlations. For classifier $f_{\beta}=\text{sign}(\beta^{*\top}x)$ with $\beta^{*}\in\mathbb{R}^{p}$, the worst-group error can be decomposed as
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\operatorname{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-\operatorname{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right).$
@@
The sign $\pm$ appearing in Theorem 1 originates from the fact that the ERM solution Under cross-entropy loss is determined only up to an overall sign. That is, both $\Sigma^{-1}\tilde{\Delta}$ and $\Sigma^{-1}\tilde{\Delta}$ define equivalent solutions, since the classifier depends only on the direction of $\beta^{*}$. Consequently, the $\pm$ in the expression of the worst-group error does not alter the actual decision boundary or error behavior, but we retain it explicitly to make the sign ambiguity transparent.
@@
Table 1: Performance comparison on multiple datasets, each exhibiting different levels of spurious correlations. For ColorMNIST, we report results under $\rho=80\%$, representing moderate spurious correlation. denotes the performance reported from yang2023change except for ColorMNIST. denotes the performance reported from deng2023robust for Waterbird and CelebA. †† denotes the performance reported from wen2025elastic for Waterbird and CelebA. For the remaining datasets not cited, detailed experimental settings following yang2023change are described in Appendix B.1.2. SCER achieves the highest worst-group accuracy in Waterbirds, CelebA, MetaShift, and ColorMNIST, demonstrating its robustness in subpopulation shift scenarios.
Algorithm Waterbirds CelebA MetaShift ColorMNIST

### v2

# arXiv:2511.04401v2[cs.LG] 02 Mar 2026

@@ Abstract
Deep learning models achieve strong performance across various domains but often rely on spurious correlations, making them vulnerable to distribution shifts. This issue is particularly severe in subpopulation shift scenarios, where models struggle in underrepresented groups. While existing methods have made progress in mitigating this issue, their performance gains are still constrained. They lack a theoretical motivation connecting the embedding space representations with worst-group error. To address this limitation, we propose Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness (SCER), a novel approach that directly regularizes feature representations to suppress spurious cues. We theoretically show that worst-group error is influenced by how strongly the classifier relies on spurious versus core directions, as identified from differences in group-wise mean embeddings across domains and classes. By imposing theoretical constraints at the embedding level, SCER encourages models to focus on core features while reducing sensitivity to spurious patterns. Through systematic evaluation on multiple vision and language tasks, we show that SCER outperforms prior state-of-the-art methods in worst-group accuracy. Our code is available at https://github.com/MLAI-Yonsei/SCER.
@@ 1 Introduction
Deep neural architectures have transformed performance standards in many domains. However, they are still highly vulnerable to spurious correlations in real-world datasets (Yang et al., 2023a; Izmailov et al., 2022; Sagawa et al., 2020). Spurious correlations are patterns in the training data that are statistically associated with target labels but lack a genuine causal relationship. Such correlations are often the result of biases in the dataset, causing models to fail when faced with distribution shifts (Ye et al., 2024). Recent studies have shown that this issue is particularly severe in subpopulation shift scenarios, where the distribution of specific subpopulations in the test data differs from that in the training data (Sagawa et al., ; Yang et al., 2023b). Models trained with Empirical Risk Minimization (ERM) tend to over-rely on these spurious correlations, as ERM minimizes the average loss without considering subpopulation imbalances (Vapnik, 2013; Izmailov et al., 2022). As a result, models rely on easily learned but non-robust features, leading to poor worst-group accuracy and unreliable performance across underrepresented subpopulations.
@@
Numerous approaches have been developed to handle this challenge, which can be broadly grouped into several categories (Yang et al., 2023b). Subpopulation robustness methods  (Sagawa et al., ; Duchi and Namkoong, 2021; Liu et al., 2021; Yao et al., 2022; Deng et al., 2023) focus on improving worst-group performance through various techniques such as reweighting, multi-stage training, or specialized loss functions. Domain invariant methods (Arjovsky et al., 2019; Sun and Saenko, 2016) attempt to learn features that generalize across different domains by aligning feature distributions. Data augmentation approaches (Zhang et al., 2018) create synthetic training examples to improve generalization. Class imbalance methods (Japkowicz, 2000; Cui et al., 2019; Ren et al., 2020) address the uneven distribution of samples across classes. However, these approaches primarily influence model learning indirectly through sample reweighting or feature alignment, without explicitly constraining how spurious features are encoded within the embedding space. Consequently, spurious correlations may persist, limiting robustness under distribution shifts.
@@
To address this limitation, we propose Spurious Correlation-Aware Embedding Regularization for Worst-Group Robustness (SCER), a novel method that directly regularizes feature representations to prevent reliance on spurious features. As illustrated in Figure 1, SCER operates at the embedding level, imposing explicit constraints that guide the model toward learning more robust features. In this study, we provide a new theoretical relationship between worst-group error and the structural properties of the representation space under subpopulation shift scenarios, where spurious correlations exist between domain and label. We identify two key factors that influence worst-group error through our theoretical analysis. First, for the same-label samples across different domains, we regularize their embeddings to align on essential classification features. Second, for instances sharing the same domain but with different labels, we introduce additional constraints to reduce the impact of domain-specific spurious cues. Through this approach, SCER prevents the model from overfitting to domain-level artifacts, enabling it to learn robust core features more effectively.
@@
By structuring the model’s representation space, SCER improves worst-group accuracy while maintaining strong overall performance. Through extensive experiments across image and text datasets, we demonstrate that SCER outperforms prior methods in worst-group accuracy. We also provide quantitative and qualitative evidence that embedding-level constraints yield robust feature representations even under spurious domain shifts, confirming the effectiveness of the SCER.
Figure 1: Overview of the Spurious Correlation-Aware Embedding Regularization (SCER) framework. Input data is encoded into embeddings, from which group-wise mean embeddings are computed for each label-domain pair. SCER decomposes these into Spurious Features, capturing domain differences within each class, and Core Features, capturing class differences within each domain. Spurious Magnitude and Core Magnitude quantify the strength of domain-driven and label-driven variations. Weight-Spurious Alignment and Weight-Core Alignment measure how strongly the classifier aligns with each component. SCER constructs Spurious Loss and Core Loss from these terms to regulate the embedding space, reducing Spurious Loss to suppress domain-specific patterns while increasing Core Loss to reinforce label-relevant signals, thereby improving robustness to domain bias.
@@ 2.1 Subpopulation Shift
Machine learning models degrade under subpopulation shift, where the test subpopulation distributions differ from those in the training data (Yang et al., 2023b). ERM’s focus on average training loss makes models susceptible to spurious correlations (Vapnik, 2013; Izmailov et al., 2022). To address this issue, various methods have been suggested, which can be broadly categorized as sample reweighting and data augmentation strategies. GroupDRO (Sagawa et al., ) reduces worst-group error by dynamically upweighting samples from groups with higher loss. LfF (Bahng et al., 2020) employs a two-stage approach that first detects spurious correlations by training a biased model and then debiases a second model by reweighting the loss gradient. GIC (Han and Zou, 2024) improves the accuracy of group inference by leveraging spurious attribute associations. For data augmentation, LISA (Yao et al., 2022) enhances robustness by interpolating between subpopulation samples to encourage invariant representations. Similarly, PDE (Deng et al., 2023) gradually expands the training set, allowing models to first learn core features before incorporating more diverse samples. While prior research focuses primarily on sample reweighting or data augmentation, we propose SCER to directly regularize the embedding space and reduce reliance on spurious features. By explicitly penalizing reliance on spurious cues at the representation level, SCER significantly improves robustness across various distribution shift benchmarks.
@@ 2.2 Embedding Space
Recent studies in the text and image domains have actively explored methods to refine latent spaces, aiming to improve representation quality and better capture the underlying structure among samples (Patil et al., 2023; Ming et al., ). However, in the domain generalization setting, research has primarily focused on learning invariant features rather than explicitly modifying the embedding space. Previous work has attempted to mitigate unintended correlations arising from background, lighting, or dataset-specific biases (Yang et al., 2023b). Hermann and Lampinen (2020) analyzed how networks differentiate robust and spurious features, while LfF  (Bahng et al., 2020) introduced a two-stage training approach to guide model learning. More recently, ElRep (Wen et al., 2025) applies norm penalties to final-layer representations with demonstrated practical effectiveness. However, these methods typically guide model learning indirectly rather than explicitly constraining spurious feature encoding in the embedding space (Patil et al., 2023; Ming et al., ). In contrast to previous approaches, we propose to directly impose an explicit loss function on the embedding space, ensuring that learned representations minimize reliance on spurious correlations and improve generalization under distribution shifts. Our work addresses these theoretical gaps by providing motivation for embedding regularization and establishing clear connections to robust generalization.
@@ 3.1 Problem Setting
To further enhance robustness, we introduce a regularization approach motivated by the decomposition of worst-group error into spurious and core components within the embedding space. Rather than addressing group errors solely through loss optimization, we incorporate structural regularization that encourages the classifier to align with label-consistent Core Feature while penalizing dependence on domain-specific Spurious Feature. These regularization terms, formulated through directional correlations in the embedding space, are detailed in the following section.
@@ 3.2 SCER
We propose SCER, which enhances the generalization performance of classifiers by decomposing the worst-group error and applying embedding-level regularization to each component, as shown in Figure 2. Let $x_{\text{emb}}=f_{w}(x)\in\mathbb{R}^{p}$ denote the embedding vectors of input $x$, obtained by a feature extractor $f_{w}$ with parameters $w$, where $p$ is the embedding dimension. For each subpopulation defined by a label–domain pair $(y,d)$, we define its mean embedding as
@@
We compute mean embeddings as representative vectors for each subpopulation, enabling decomposition of worst-group error into two primary components.
Spurious Feature. The embedding mean difference in feature representations across different domains within the same class. Since we assume that label–domain pairs $(y,d)$ may exhibit spurious correlations, such differences should be minimized to prevent the model from overfitting to domain-specific artifacts. Formally, the spurious difference for a given class $y$ is defined as
@@
Core Feature. The embedding mean difference in feature representations between different classes within the same domain. This quantity captures the actual signal that the classifier should leverage for robust prediction. Formally, the core difference for a given domain $d$ is defined as
@@
To ensure proper interpretation based on theoretical insights, we normalize these direction vectors using the $\Sigma$-norm, defined as $\|v\|_{\Sigma}=\sqrt{v^{\top}\Sigma v}$ for $v\in\mathbb{R}^{p}$, where $\Sigma\in\mathbb{R}^{p\times p}$ denotes the empirical covariance matrix of the embedding vectors. This normalization accounts for the geometric structure of the embedding space, producing scalar quantities Spurious Magnitude $\|\Delta_{\text{spur}}\|_{\Sigma}$ and Core Magnitude $\|\Delta_{\text{core}}\|_{\Sigma}$ for use in regularization. We also derive regularization terms by analyzing how classifier weights align with the Spurious and Core Directions. The key idea is that overreliance on Spurious Directions leads to poor generalization on minority groups, while alignment with Core Directions supports robust performance. Let $\beta^{*}=[\beta^{*}_{1},\dots,\beta^{*}_{j},\dots,\beta^{*}_{m}]\in\mathbb{R}^{p\times m}$ denote the current weight matrix of the classifier $f_{\beta}$, where each column $\beta^{*}_{j}\in\mathbb{R}^{p}$ corresponds to the weight vector for class $j$. During training, we use the current classifier weights to compute the alignment measures. To quantify the alignment between classifier weights and directional signals, we compute the average class-wise correlation with each direction using the $\Sigma$-norm. We refer to these measures as Weight-Spurious Alignment and Weight-Core Alignment
@@
These correlation terms measure how closely the classifier’s decision boundaries align with spurious or core feature directions, providing the basis for our embedding-level regularization.
@@
The Spurious Loss $\mathcal{L}_{\text{spur}}$ penalizes alignment with spurious directions, scaled by the magnitude of spurious variation in the data. The Core Loss $\mathcal{L}_{\text{core}}$ encourages alignment with core directions, with the negative sign ensuring that more substantial core alignment reduces the overall loss. Our loss formulation penalizes alignment with spurious directions and encourages alignment with core directions. By introducing control parameters $\lambda_{\text{spur}}$ and $\lambda_{\text{core}}$, we derive our embedding loss function:
@@ 3.3 Theoretical Analysis of SCER
We now present a theoretical analysis that formally decomposes worst-group error. The following theorem demonstrates that worst-group error can be decomposed into the classifier’s reliance on spurious versus core components, thereby providing theoretical motivation for SCER’s regularization of these components. Detailed proofs are provided in Appendix A.2.
@@
Consider a classification problem where each embedding vector $x\in\mathbb{R}^{p}$ is associated with a label $\mathcal{Y}=\{y_{-1},y_{+1}\}$ and a domain $\mathcal{D}=\{d_{R},d_{G}\}$. Each pair $(y,d)\in\mathcal{Y}\times\mathcal{D}$ defines a subpopulation. We assume that the data follows a group-conditional Gaussian distribution $x\mid(y,d)\sim\mathcal{N}(\mu^{(y,d)},\Sigma)$ where $\Sigma$ is positive definite. We further assume that $\Delta_{\text{core}}=\mu^{(y_{+1},d)}-\mu^{(y_{-1},d)}$ is constant across $d$, $\Delta_{\text{spur}}=\mu^{(y,d_{R})}-\mu^{(y,d_{G})}$ is constant across $y$, $\Sigma^{(d_{R})}=\Sigma^{(d_{G})}=\Sigma$, $\mathbb{P}(y=\pm 1)=\mathbb{P}(d\in\{d_{R},d_{G}\})=\frac{1}{2}$, and $\mathbb{P}(y,d)\approx 0$ for some pairs $(y,d)$ under extreme spurious correlations. For classifier $f_{\beta}=\text{sign}(\beta^{*\top}x)$ with $\beta^{*}\in\mathbb{R}^{p}$, the worst-group error can be decomposed as
$E_{\mathrm{wge}}=\Phi\!\left(\pm\tfrac{1}{2}\,\operatorname{cor}(\beta^{*},\Delta_{\mathrm{spur}})\,\|\Delta_{\mathrm{spur}}\|_{\Sigma}-{\tfrac{1}{2}\,}\operatorname{cor}(\beta^{*},\Delta_{\mathrm{core}})\,\|\Delta_{\mathrm{core}}\|_{\Sigma}\right).$
@@
The sign $\pm$ in Theorem 1 reflects which subgroup constitutes the worst group. When $\Delta_{\mathrm{spur}}^{\top}\beta^{*}>0$, the unseen minority subgroups incur higher error, corresponding to the ++ sign; the - sign arises in the rare case where seen subgroups are worst. In the typical ERM regime with spurious correlations, the ++ sign applies, and we retain $\pm$ for generality. See Appendix A.2 for the detailed derivation.
@@
Table 1: Performance comparison on multiple datasets, each exhibiting different levels of spurious correlations. denotes the performance reported from Yang et al. (2023b) except for ColorMNIST. denotes the performance reported from Deng et al. (2023) for Waterbird and CelebA. †† denotes the performance reported from Wen et al. (2025) for Waterbird and CelebA. For the remaining datasets, the detailed experimental settings follow Yang et al. (2023b), as described in Appendix B.1.2. SCER achieves the highest worst-group accuracy in Waterbirds, CelebA, MetaShift, and ColorMNIST, demonstrating its robustness in subpopulation shift scenarios.
Algorithm Waterbirds CelebA MetaShift ColorMNIST ($\rho=80\%$)